Combustion reactions are a notable source of carbon dioxide in the environment. Using the following balanced equation, how many grams of carbon dioxide are formed when 100.00 g of propane (C3H8) is burned? Express your answer to the correct number of significant figures. Be sure to show all steps completed to arrive at the answer.

C3H8 + 5O2 = 3CO2 + 4 H2O


1
Expert's answer
2021-06-18T07:23:24-0400

C3H8 + O2 = 3CO2 + 4 H2O

Moles of C3H8 = 100 / 44 = 2.28 mole

Moles of CO2 = ( 3 × 2.28 ) mole.

Mass of CO2 = ( 3× 2.28 )× 44 gram

= 300.96 gram .